# Question-to-Question Retrieval for Hallucination-Free Knowledge Access: An Approach for Wikipedia and Wikidata Question Answering

### v2

@@ 1 Introduction
This paper introduces a novel approach to knowledge-based question answering that addresses these limitations by performing "question-to-question" retrieval and avoids the generation step. Instead of embedding document content, we employ an instruction-tuned LLM to generate a comprehensive set of questions for each logical content unit within the knowledge base. These generated questions are then vector-embedded and stored in a searchable vector store, mapped to the corresponding source document. When a user submits a query, it is also embedded in the same space and a search performed to identify the most similar generated question. For highest matching question, we present the corresponding article and the paragraph with in the article. A similar approach is presented for wikidata where the content is in the form of enitity relations, we generate all possible questions and then make the wikidata knowledge graph accessible to a question answer system.
@@ 2 Background
Assume that our collection contains $D$ documents, $d_{1},d_{2},\cdots,d_{D}$. We first split each of the Wikipedia article into text passages as the basic retrieval units and get $M$ total passages in our corpus $C={p_{1},p_{2},\cdots,p_{M}}$, where each passage $p_{i}$ can be viewed as a sequence of tokens $w^{(i)}_{1},w^{(i)}_{2},\cdots,w^{(i)}_{|p_{i}|}$ Given a question $q$, the task is to find a $p$ that can answer the question. For the sake of quick and precice mapping between passage and article, assume there is a unique cryptographic hash based on the content of each $p$ that is mapped to a $d$ and saved in a relational database.
@@ 3 Methodology
The Wikipedia content is primarily in Wikitext markup. It can also be rendered to HTML. But for the purpose of our system, we prepared plain text version of earch passage so that it is comprehensible for an LLM. We also removed reference numbers that appear in usual plain text format.
@@
For each content unit $d_{i}\in\mathcal{D}$, a set of questions is generated by an LLM, denoted as $Q_{i}=\{q_{i1},q_{i2},...,q_{in_{i}}\}$, where $n_{i}$ is the number of questions generated for $d_{i}$. The prompt used for the LLM is given in Appendix LABEL:sec:llm-prompt-wikipedia-questions. The input to the prompt not only contains the passage text, but contextual information like article title, section titles to resolve coreferences. The prompt also uses few shot prompting to get machine readable output.
@@ 3.1 QA on Wikidata
To incorporate Wikidata, we needed to extract factual information from the knowledge base in a format that is suitable for the question generation task. We wrote a SPARQL query(See Listing 1) that retrieves all subjects, predicates, and objects (triples) associated with each Wikidata entity (QID). The query is designed to be comprehensive in extracting fact triples in a human-readable textual format (e.g., “India: inception: 15 August 1947” or "India: Prime Minister: Narendra Modi (2014-current)"). A unique hash of QID and PID pair is calculated for linking the extracted content with generated questions.

### v3

@@ 1 Introduction
This paper introduces a novel approach to knowledge-based question answering that addresses these limitations by performing "question-to-question" retrieval and avoids the generation step. Instead of embedding document content, we employ an instruction-tuned LLM to generate a comprehensive set of questions for each logical content unit within the knowledge base. These generated questions are then vector-embedded and stored in a searchable vector store, mapped to the corresponding source document. When a user submits a query, it is also embedded in the same space and a search performed to identify the most similar generated question. For highest matching question, we present the corresponding article and the paragraph with in the article. A similar approach is presented for wikidata where the content is in the form of entity relations, we generate all possible questions and then make the wikidata knowledge graph accessible to a question answer system.
@@ 2 Background
Assume that our collection contains $D$ documents, $d_{1},d_{2},\cdots,d_{D}$. We first split each of the Wikipedia article into text passages as the basic retrieval units and get $M$ total passages in our corpus $C={p_{1},p_{2},\cdots,p_{M}}$, where each passage $p_{i}$ can be viewed as a sequence of tokens $w^{(i)}_{1},w^{(i)}_{2},\cdots,w^{(i)}_{|p_{i}|}$ Given a question $q$, the task is to find a $p$ that can answer the question. For the sake of quick and precise mapping between passage and article, assume there is a unique cryptographic hash based on the content of each $p$ that is mapped to a $d$ and saved in a relational database.
@@ 3 Methodology
The Wikipedia content is primarily in Wikitext markup. It can also be rendered to HTML. But for the purpose of our system, we prepared plain text version of each passage so that it is comprehensible for an LLM. We also removed reference numbers that appear in usual plain text format.
@@
For each content unit $d_{i}\in\mathcal{D}$, a set of questions is generated by an LLM, denoted as $Q_{i}=\{q_{i1},q_{i2},...,q_{in_{i}}\}$, where $n_{i}$ is the number of questions generated for $d_{i}$. The prompt used for the LLM is given in Appendix LABEL:sec:llm-prompt-wikipedia-questions. The input to the prompt not only contains the passage text, but contextual information like article title, section titles to resolve coreferences. The prompt also uses few shot prompting to get machine readable output. The concept of document expansion or enrichment by adding queries is common in information retrieval techniquesNogueira et al. (2019). Here we are not enhancing the document(passage), but uses the questions generated out of it.
@@ 3.1 QA on Wikidata
To incorporate Wikidata, we needed to extract factual information from the knowledge base in a format that is suitable for the question generation task. We wrote a SPARQL query(See Listing 1) that retrieves all subjects, predicates, and objects (triples) associated with each Wikidata entity (QID). The query is designed to be comprehensive in extracting fact triples in a human-readable textual format (e.g., “India: inception: 15 August 1947” or "India: Prime Minister: Narendra Modi (2014-current)"). The conversion of triples to text for querying on structured content was originally discussed in UniK-QA systemOguz et al. (2022). A unique hash of QID and PID pair is calculated for linking the extracted content with generated questions.
@@ 5 Conclusion
Acknowledgments
This paper benefited from the valuable contributions of Isaac Johnson and Xiao Xiao of the Wikimedia Foundation.